The perimeter of a rectangle is 60 units. Its width is 5 units. What is the area of the rectangle?

Answer by mananth(16946)

width is 5 units
2 times width = 10
2L+2W = 60
60-2w=2l
60-10=50
50 = 2 length
length = 25 units
Area = 25* 5 = 125 sq. units